# ASR-enhanced Multimodal Representation Learning for Cross-Domain Product Retrieval

### v1

@@ Abstract
Multimodal Analysis, ASR Text Summarization, Multimodal Product Representation, Cross-domain Product Retrieval.
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ I Introduction
The content of E-commerce is becoming increasingly multimedia-enriched. Besides traditional product pages, E-commerce platforms are incorporating more dynamic and interactive media formats, such as short videos and live streams, to present their products more effectively and consequently enhance end-user shopping experience. A specific product can now be exhibited in a broad-domain manner through an image [1], a short descriptive video [2], or a live stream promotion [3], see Fig. 1. As such, a unified and vectorized cross-domain representation of the product is essential for downstream tasks including cross-domain product retrieval (CdPR), user profiling, personalized recommendation, etc. This paper is targeted at multimodal product representation learning (MmPRL) for CdPR.
@@
To attack the challenge, Bai et al. develop a tri-domain benchmark termed ROPE [3], containing millions of multimodal samples from three domains, i.e. product-page, short-video and live-stream. Their proposed solution for CdPR is fully reliant on the visual content, which alone is insufficient to handle large intra-product variance and high inter-product similarity, see Fig. 1. Chen et al. propose a query-driven approach [2], assuming that a product image or a short video is associated with a textual query based on click-through records. However, such a text modality is often unavailable in particular for novel products. In a contemporary study, Yang et al. propose to exploit Automatic Speech Recognition (ASR) text transcribed from the audio channel of short / live-stream videos [13]. Despite its wide availability, the ASR text is typically overwhelmed by abundant uninformative words from casual chatting of network anchors or live streamers, see Table I. Indeed, the raw ASR text is previously deemed to be harmful for product representation [3].
TABLE I: Our prompt to let an LLM (Baichuan2-13B-Chat-4bits) perform ASR text summarization. We experiment with Chinese data, so the prompt used is written in Chinese. We show its English counterpart for non-Chinese readers.
@@
Although ASR text has been used in varied video tasks such as text-video retrieval [14], video representation learning [15], video abstractive summarization [16], and video question answering [17], how to de-noise the text for multimodal representation learning is almost untouched. Given the excessive noise in ASR text derived from E-commence videos, our experiment shows that an advanced information extraction tool as UIE [18] fails to extract anything for over 30% of the ROPE data. Inspired by the great success of large language models (LLMs), we propose to de-noise the ASR text with an LLM-based text summarizer, see Table I. MmPRL with the enhanced ASR text leads to much improved cross-domain product retrieval results, as demonstrated in Fig. 1. A conceptual diagram of our solution is illustrated in Fig. 2. In short, our major contributions are as follows:
  • We propose ASR-enhanced Multimodal Product Representation Learning (AMPere), showing for the first time that incorporating ASR text can largely benefit product retrieval in the E-commerce domain.
@@
Figure 2: Conceptual diagram of the proposed AMPere method for cross-domain multimodal product representation learning. Our network consists of three branches, each responding to a specific domain. For each trainable layer in the network, its parameters are shared across the three branches. An LLM-based ASR text summarizer is deployed to extract product-specific information from the raw text overwhelmed by abundant uninformative words transcribed from casual chatting of network anchors or live streamers. Except for the summarizer, the network is trained end-to-end.
@@ II Related Work
This work is targeted at learning ASR-enhanced multimodal production representation for cross-domain product retrieval (CdPR). Hence, we first discuss recent progress in CdPR, followed by a brief survey concerning the use of ASR text in a broader context.
@@ II-A Cross-domain Product Retrieval
Compared to single-domain (mostly product-page) product retrieval [5, 6, 7, 9, 8, 10, 11, 19, 20], there are relatively few studies on CdPR, see Table II. Chen et al. [2] propose QCD for bidirectional retrieval between the product-page (P) and short-video (S) domains. In particular, QCD assumes each visual example, either a product image in the P domain or a short video in the S domain, is associated with a user query obtained based on click-through records. By cross-attention based fusion of the textual feature of the query and the visual feature extracted from the visual example, a multimodal product representation (MmPR) is obtained. Since user queries are relatively short, no text enhancement is involved in QCD. Moreover, for videos lacking click-through in particular those related to newly listed products, no query will be available. In [13], Yang et al. propose RICE for retrieving product pages for a given live-stream video. In particular, given a product image with a title and a video with ASR text, their cross-domain similarity is calculated by combining the video-image similarity and the ASR-title similarity. No ASR text enhancement is conducted. In a contemporary work [3], Bai et al. consider a more realistic yet more challenging scenario of three-domain CdPR. However, their proposed COPE method is fully visual based, without exploiting the text modality [^1] After analyzing the official code of COPE, we find that the title channel of the product-page domain was also unused.. We follow the challenging setting of COPE, and go one step further by exploiting the ASR text for obtaining MmPR.
@@
QCD, MM’23[2] P, S P2S, S2P user query
[…]
2613
RICE, ICCV’23 [13] P, L L2P ASR text
[…]
2613
[…]
2613
COPE, ICCV’23 [3] P, S, L P2L, P2S, S2P, S2L, L2P, L2S
[…]
2613
[…]
2613
AMPere P, S, L P2L, P2S, S2P, S2L, L2P, L2S ASR text
@@ II-B The Use of ASR Text for Video Analysis
The ASR text has been frequently used in varied video tasks. In text-video retrieval, for instance, the ASR text is typically vectorized by word2vec and used together with other features [14, 21, 22]. The text is also used as relatively weak annotations for large-scale pre-training [23, 24, 25]. All the above works use the ASR text as is albeit its noisy nature. For VideoQA, Zellers et al. [17] propose to add punctuation to the ASR with a sequence-to-sequence model that is trained to add punctuation to sentences / paragraphs from news articles. For generic text-video retrieval, Chen et al. [26] employ KeyBERT [27] to extract keywords from the ASR text. With the assumption that words most similar to a given document are good keywords for representing the document, KeyBERT calculates word-document similarities based on their BERT embeddings. In the E-commerce field, live streamers often spend much of their time chatting casually during product promotion sessions. The actual time dedicated to discussing the products is typically a small portion. As such, the ASR text transcribed from short or live-stream videos is too noisy to be handled by KeyBERT. To conquer the challenge, we make a novel attempt by harnessing a large language model (LLM) for effective ASR text summarization.
@@ III-A Problem Formalization
We formalize the problem of multimodal product representation learning (MmPRL) as follows. A specific product on an E-commerce platform is typically exhibited as data samples in the following three domains, i.e. product-page (P), short-video (S), and live-stream (L). Let $x$ be a specific sample, with $x_{p}$, $x_{s}$ and $x_{l}$ be its three domain-specific instantiations. In particular, $x_{p}$ indicates an image of the given product, $x_{s}$ is a short video describing the product, whilst $x_{l}$ denotes a longer live-stream video with a network anchor prompting the product. In a multimodal scenario, the image $x_{p}$ is associated with a title $t_{p}$, while the two videos $x_{s}$ and $x_{l}$ are associated with ASR texts $t_{s}$ and $t_{l}$, respectively. For the ease of consistent description, we re-use $x$ to indicate a given multimodal sample, with its domain identity and the associated text both temporally omitted. The goal of MmPRL is to train a feature extraction network $\mathcal{F}$ that encodes the given sample into a $d$-dimensional embedding $e(x)$. Accordingly, cross-domain product retrieval boils down to calculating cosine similarities between $e(x_{p})$, $e(x_{s})$ and $e(x_{l})$.
@@ III-B COPE in a Nutshell
$\left.\begin{array}[]{ll}\{f_{1},\ldots,f_{n}\}&\leftarrow\mbox{video-to-frames}(x,n=8),\\ \{z_{1},\ldots,z_{n}\}&\leftarrow\mbox{CCT}(\{f_{1},\ldots,f_{n}\}),\\ v(x)&\leftarrow\mbox{AvgPool}(\mbox{Transformers}(\{z_{1},\ldots,z_{n}\})),\\ e(x)&\leftarrow\mbox{Linear}_{512\times 128}(v(x)).\end{array}\right.$ (1)
@@ III-C LLM-based ASR Text Summarization
Given the verbose ASR text obtained from a video of a specific product, say an electric shaver, we aim to extract two key pieces of product-specific information. That is, the name of the product (e.g. Space UFO shaped shaver) and major features of the product (e.g. high quality blade material, fast and clean shaving, …). To that end, we turn an LLM into an ASR text summarizer with the popular In-Context Learning (ICL) technique [30]. In essence, ICT performs a test-time adaptation of an LLM by instructing the model with a few demonstration examples. See Table I for our demonstrations and our project website [^2] https://rucmm.github.io/mmrl4cdpr/ for source code.
As our benchmark dataset [3] was gathered from a Chinese E-commerce platform with text in Chinese, a Chinese LLM is needed. In the current work we adopt a 4bits quantized version of the 13B-parameter Baichuan Chat model [^3] https://huggingface.co/baichuan-inc/Baichuan2-13B-Chat, a leading LLM on the CLiB Chinese LLM benchmark [^4] https://github.com/jeinlee1991/chinese-llm-benchmark. This version strikes a good balance between model effectiveness and inference cost. Qualitative results of our LLM-based ASR text summarization are provided in Fig. 3.
@@
Figure 3: Qualitative results of ASR text summarization. We compare with KeyBERT [27], a BERT embedding based keyword extraction tool, and UIE [18], a competitive open-source model for information extraction. Note the text is shown partially due to space limit. The attention map per frame is obtained with the CLS token as queries (Q) and the patch tokens as keys (K). Since the CLS token is used as the frame-level feature for subsequent video feature extraction, the attention map shows the region the model is focusing on, for instance, the small hot spots corresponding to earrings in Fig. 3a. Best viewed digitally.
@@ III-D Adding Summarized ASR to COPE
For each trainable layer in the network, its parameters are shared across the three branches. We empirically find that such a simple parameter-sharing strategy is better than alternatives such as training the last linear layer in a domain-specific manner.
Compared to COPE which has 134.4M parameters, AMPere has 206.8M parameters due to the use of the text encoder and the fusion modules. Except for the ASR text summarizer, our network is trained in an end-to-end manner by minimizing the same loss as COPE. Once trained, the network is used to produce multimodal embeddings $e(x_{p},t_{p})$, $e(x_{s},t_{s})$, $e(x_{l},t_{l})$ for bidirectional CdPR between the three domains.
@@ IV-A Experimental Setup
Dataset. We adopt ROPE [3], a public dataset released by Kuaishou (Kwai) for large-scale cross-domain product retrieval [^5] https://github.com/adxcreative/COPE. With 3.1M product pages, 5.9M short videos, and 3.5M live streams for over 189K products, ROPE is currently the largest multi-domain multimodal product dataset of its kind. Tab. III shows data statistics of ROPE. We follow the official data split, 98.7% for training and 1.3% for testing.
TABLE III: Statistics of the ROPE dataset [3] used in this study. Text length is the number of Chinese characters per sample. All values are means except for #Images and #Videos.
@@
Product page #Images 3.0M 31.4K
Text length $35.0$ $21.5$
Short video #Videos 5.8M 29.8K
Duration (s) $35.4$ $31.8$
Text length:
@@
Live stream #Videos 3.5M 31.0K
Duration (s) $112.2$ $129.09$
Text length:
@@
Evaluation criteria. Per task, we report standard rank-based metrics: R1, R5 and R10. For overall comparison, we report the mean value of R1 of all tasks (mR1).
Implementation details. As mentioned earlier, our solution is developed on top of COPE [3]. So for fair comparison, we follow the implementation of this baseline wherever applicable: the video encoder is initialized by X-CLIP-B/16 [28] pretrained on Kinetics-600 [32], the number of frames $n$ is set to 8, and a combined loss of an instance-level contrastive loss [33] and a product-level classification loss [34] is minimized by the AdamW optimizer [35]. The maximum token length for the text encoder is empirically set to 32. The batch size is 128. The learning rate is first warmed up by 4 epochs to 5e-5, 1e-5, 5e-5, and 5e-3 for the text encoder, the video encoder, the fusion transformers, and the other layers, respectively, and then follows a cosine annealing decay schedule. For all models, we obtain their stable checkpoints at the 70th training epoch. Experiments are conducted with PyTorch on 16 NVIDIA Tesla V100 GPUs.

### v2

@@ Abstract
ASR Text Summarization, Multimodal Product Representation, Cross-domain Product Retrieval
publicationid: pubid: $0$–$00$ 2025 IEEE
@@ I Introduction
The content of E-commerce is becoming increasingly multimedia-enriched. Besides traditional product pages, E-commerce platforms are incorporating more dynamic and interactive media formats, such as short videos and live streams, to present their products more effectively and consequently enhance end-user shopping experience. A specific product can now be exhibited in a broad-domain manner through an image [1], a short descriptive video [2], or a live stream promotion [3], see Fig. 1. As such, a unified and vectorized cross-domain representation of the product is essential for downstream tasks including cross-domain product retrieval (CdPR) [3], user profiling, personalized recommendation, and online video advertising [4, 5], to name just a few. This paper is targeted at multimodal product representation learning (MmPRL) for CdPR.
@@
To attack the challenge, Bai et al. develop a tri-domain benchmark termed ROPE [^1] https://github.com/adxcreative/COPE, containing millions of multimodal samples from three domains, i.e. product-page, short-video and live-stream [3]. Their proposed solution for CdPR is fully reliant on the visual content, which alone is insufficient to handle large intra-product variance and high inter-product similarity, see Fig. 1. Chen et al. propose a query-driven approach [2], assuming that a product image or a short video is associated with a textual query based on click-through records. However, such a text modality is often unavailable in particular for novel products. In a contemporary study, Yang et al. propose to exploit Automatic Speech Recognition (ASR) text transcribed from the audio channel of short / live-stream videos [15]. Despite its wide availability, the ASR text is typically overwhelmed by abundant uninformative words from casual chatting of network anchors or live streamers, see Table I. Indeed, the raw ASR text is previously deemed to be harmful for product representation [3].
TABLE I: Our prompt to let an LLM (Baichuan2-13B-Chat-4bits) perform ASR text summarization. We experiment with Chinese data, so the prompt used is written in Chinese. Its English counterpart is shown for non-Chinese readers.
@@
Although ASR text has been used in varied video tasks such as text-video retrieval [16], video representation learning [17, 18], video abstractive summarization [19],video question answering [20] and video hyperlinking [21, 22], how to de-noise the text for multimodal representation learning is almost untouched. Given the excessive noise in ASR text derived from E-commence videos, we empirically find that an advanced information extraction tool as UIE [23] fails to extract anything for over 30% of the ROPE data. Inspired by the great success of large language models (LLMs), we propose to de-noise the ASR text with an LLM-based text summarizer, see Tab. I. MmPRL with the enhanced ASR text leads to much improved CdPR results. In short, our major contributions are as follows:
  • We propose ASR-enhanced Multimodal Product Representation Learning (AMPere), see Fig. 2, showing for the first time that incorporating ASR text can largely benefit product retrieval in the E-commerce domain.
@@
Figure 2: Proposed AMPere method for cross-domain multimodal product representation learning. Our network consists of three branches, each responding to a specific domain, i.e. short-video, live-stream and product-page. For each trainable layer in the network, its parameters are shared across the three branches. An LLM-based ASR text summarizer is deployed to extract product-specific information from the raw text overwhelmed by abundant uninformative words transcribed from casual chatting of network anchors or live streamers. Except for the summarizer, the network is trained end-to-end.
@@ II Related Work
This work is targeted at learning ASR-enhanced multimodal production representation for CdPR. Hence, we first discuss recent progress in CdPR, followed by a brief survey concerning the use of ASR text in a broader context.
@@ II-A Cross-domain Product Retrieval
Despite good attempts on cross-domain image-to-image retrieval in a non-ecommerce context [24] and single-domain (mostly product-page) product retrieval [7, 8, 9, 11, 10, 12, 13, 25, 26], there are relatively few studies on cross-domain product retrieval, see Table II. In an earlier work [27], Cheng et al. make a novel attempt to match clothes appeared in videos to the same items in online shops by cross-domain visual feature matching. In particular, a clothing detector is used to localize regions of interest in a given video. Given the difficulty in training good detectors for a wide range of products, their detector-based solution is not directly applicable in the current context that has tens of thousands of diverse products. Moreover, as [27] is purely content-based, no multimodal representation learning is involved. In [2], Chen et al. propose QCD for bidirectional retrieval between the product-page (P) and short-video (S) domains. In particular, QCD assumes each visual example, either a product image in the P domain or a short video in the S domain, is associated with a user query obtained based on click-through records. By cross-attention based fusion of the textual feature of the query and the visual feature extracted from the visual example, a multimodal product representation (MmPR) is obtained. Since user queries are relatively short, no text enhancement is involved in QCD. Moreover, for videos lacking click-through in particular those related to newly listed products, no query will be available. In [15], Yang et al. propose RICE for retrieving product pages for a given live-stream video. In particular, given a product image with a title and a video with ASR text, their cross-domain similarity is calculated by combining the video-image similarity and the ASR-title similarity. No ASR text enhancement is conducted. In a contemporary work [3], Bai et al. consider a more realistic yet more challenging scenario of three-domain CdPR. However, their proposed COPE method is fully visual based, without exploiting the text modality [^2] After analyzing the official code of COPE, we find that the title channel of the product-page domain was also unused.. We follow the challenging setting of COPE, and go one step further by exploiting the ASR text for obtaining MmPR.
@@
Video2Shop, CVPR17[27] P, S S2P
QCD, MM23[2] P, S P2S, S2P user query
RICE, ICCV23[15] P, L L2P ASR text
COPE, ICCV23[3] P, S, L P2L, P2S, S2P, S2L, L2P, L2S
AMPere (this paper) P, S, L P2L, P2S, S2P, S2L, L2P, L2S ASR text
@@ II-B The Use of ASR Text for Video Analysis
The ASR text has been frequently used in varied video tasks. In text-video retrieval, for instance, the ASR text is typically vectorized by word2vec and used together with other features [16, 28, 29]. For video hyperlinking, [21, 22] extract bag-of-words feature from the ASR text. For video decoration with sound effects, the ASR text is jointly exploited with video frames for multimodal video embedding [18]. The text is also used as relatively weak annotations for large-scale pre-training [30, 31, 17]. All the above works use the ASR text as is albeit its noisy nature. For VideoQA, Zellers et al. [20] propose to add punctuation to the ASR with a sequence-to-sequence model that is trained to add punctuation to sentences / paragraphs from news articles. For generic text-video retrieval, Chen et al. [32] employ KeyBERT [33] to extract keywords from the ASR text. With the assumption that words most similar to a given document are good keywords for representing the document, KeyBERT calculates word-document similarities based on their BERT embeddings. In the E-commerce field, live streamers often spend much of their time chatting casually during product promotion sessions. The actual time dedicated to discussing the products is typically a small portion. As such, the ASR text transcribed from short or live-stream videos is too noisy to be handled by KeyBERT. To conquer the challenge, we make a novel attempt by harnessing an LLM for effective ASR text summarization.
@@ III-A Problem Formalization
We formalize the problem of multimodal product representation learning (MmPRL) as follows. A specific product on an E-commerce platform is typically exhibited as data samples in the following three domains, i.e. product-page (P), short-video (S), and live-stream (L). Let $x$ be a specific sample, with $x_{p}$, $x_{s}$ and $x_{l}$ be its three domain-specific instantiations. In particular, $x_{p}$ indicates an image of the given product, $x_{s}$ is a short video describing the product, whilst $x_{l}$ denotes a longer live-stream video with a network anchor prompting the product. In a multimodal scenario, the image $x_{p}$ is associated with a title $t_{p}$, while the two videos $x_{s}$ and $x_{l}$ are associated with ASR texts $t_{s}$ and $t_{l}$, respectively. For the ease of consistent description, we re-use $x$ to indicate a given multimodal sample, with its domain identity and the associated text both temporally omitted. The goal of MmPRL is to train a feature extraction network $\mathcal{F}$ that encodes the given sample into a $d$-dimensional embedding $e(x)$. Accordingly, CdPR boils down to calculating cosine similarities between $e(x_{p})$, $e(x_{s})$ and $e(x_{l})$.
@@ III-B COPE in a Nutshell
$\left.\begin{array}[]{ll}\{f_{1},\ldots,f_{n}\}&\leftarrow\mbox{video-to-frames}(x,n=8),\\ \{z_{1},\ldots,z_{n}\}&\leftarrow\mbox{CCT}(\{f_{1},\ldots,f_{n}\}),\\ v(x)&\leftarrow\mbox{AvgPool}(\mbox{Transformers}(\{z_{1},\ldots,z_{n}\})),\\ e(x)&\leftarrow\mbox{Linear}_{512\times 128}(v(x)).\end{array}\right.$
@@ III-C LLM-based ASR Text Summarization
Given the verbose ASR text obtained from a video of a specific product, say an electric shaver, we aim to extract two key pieces of product-specific information. That is, the name of the product (e.g. Space UFO shaped shaver) and major features of the product (e.g. high quality blade material, fast and clean shaving, …). To that end, we turn a generic LLM into an ASR text summarizer with the popular In-Context Learning (ICL) technique [36], which requires no model re-parameterization nor fine-tuning. In essence, ICT performs a test-time adaptation of an LLM by instructing the model with a few demonstration examples. See Table I for our demonstrations and our project website [^3] https://rucmm.github.io/mmrl4cdpr/ for source code.
As the ROPE dataset [3] used in this study was gathered from a Chinese E-commerce platform with text in Chinese, a Chinese LLM is needed. Specifically, we adopt a 4bits-quantized version of the 13B-parameter Baichuan Chat model [^4] https://huggingface.co/baichuan-inc/Baichuan2-13B-Chat, a leading LLM on the CLiB Chinese LLM benchmark [^5] https://github.com/jeinlee1991/chinese-llm-benchmark. This version strikes a good balance between model effectiveness and inference cost. See Fig. 3 for qualitative results of our LLM-based ASR text summarization.
[…]
@@
Figure 3: Qualitative results of ASR text summarization. We compare with KeyBERT [33], a BERT embedding based keyword extraction tool, and UIE [23], a competitive open-source model for information extraction. Text is shown partially due to space limit. The attention map per frame is obtained with the CLS token as queries (Q) and the patch tokens as keys (K). Since the CLS token is used as the frame-level feature for subsequent video feature extraction, the attention map shows the region the model is focusing on, for instance, the small hot spots corresponding to earrings in Fig. 3a. Best viewed digitally.
@@ III-D Adding Summarized ASR to COPE
Due to the use of the text encoder and the fusion modules, AMPere has more parameters than COPE (207M vs. 134M). By executing Eq. 2 per branch, we obtain multimodal embeddings $e(x_{p},t_{p})$, $e(x_{s},t_{s})$, $e(x_{l},t_{l})$ for the P, S and L domains, respectively. Next, we describe how the three branches are jointly trained to ensure cross-domain embedding alignment.
III-E Cross-domain Embedding Alignment
Following COPE, we train our three-branch network by minimizing a combined loss which consists of three inter-domain pairwise contrastive losses and three intra-domain pointwise classification losses. In particular, given two specific domains, say P and S, an inter-domain contrastive loss is computed by treating $e(x_{p},t_{p})$ and $e(x_{s},t_{s})$ of the same product as a positive pair, with their counterparts from different products as negative pairs. Meanwhile, for each domain a classification loss is calculated by predicting the product label of the current instance based on its embedding $e(x,t)$. By jointly optimizing these six losses, multimodal embeddings of instances from different domains are product-wise aligned.
For each trainable layer in the network, its parameters are shared across the three branches, i.e. a single set of model parameters is used. We empirically find that such a simple parameter-sharing strategy is better than alternatives such as training the last linear layer in a domain-specific manner. Except for the LLM, our network is end-to-end trained. Once trained, the network is used to produce multimodal embeddings $e(x_{p},t_{p})$, $e(x_{s},t_{s})$, $e(x_{l},t_{l})$ for bidirectional CdPR between the three domains.
@@ IV-A Experimental Setup
Dataset. We adopt ROPE [3], a public dataset released by Kuaishou (Kwai) for large-scale CdPR. With 3.1M product pages, 5.9M short videos, and 3.5M live streams for over 189K products, ROPE is currently the largest multi-domain multimodal product dataset of its kind. Tab. III shows data statistics of ROPE. We follow the official data split, 98.7% for training and 1.3% for testing.
TABLE III: Statistics of the ROPE dataset we use. Text length is the amount of Chinese characters per sample.
@@
Product-page #Images 3.0M 31.4K
Average text length $35.0$ $21.5$
Short-video #Videos 5.8M 29.8K
Average duration $35.4$s $31.8$s
ASR missing 0.5% 1.2%
LLM No-output 9.9% 25.1%
Average text length:
@@
Live-stream #Videos 3.5M 31.0K
Average duration $112.2$s $129.1$s
ASR missing 0.4% 0.1%
LLM No-output 3.6% 3.3%
Average text length:
@@
Evaluation criteria. Per task, we report standard rank-based metrics: R1, R5 and R10. i.e. recall at top k (k=1, 5, 10). For overall comparison, we use mean R1 of all tasks (mR1). In addition, we report Mean Reciporal Rank (MRR) and Normalized Discounted Cumulative Gain at 10 (NDCG10).
Implementation details. As mentioned earlier, our solution is developed on top of COPE [3]. So for fair comparison, we follow the implementation of this baseline wherever applicable: the video encoder is initialized by X-CLIP-B/16 [34] pretrained on Kinetics-600 [38], the number of frames $n$ is set to 8. The instance-level contrastive loss [39] and the product-level classification loss [40] is jointly minimized by the AdamW optimizer [41]. The maximum length of text tokens is empirically set to 32. The batch size is 128. The learning rate is first warmed up by 4 epochs to 5e-5, 1e-5, 5e-5, and 5e-3 for the text encoder, the video encoder, the fusion transformers, and the other layers, respectively, and then adjusted with a cosine annealing decay schedule. For all models, we obtain their stable checkpoints at the 70th training epoch. Experiments are conducted with PyTorch on 16 NVIDIA Tesla V100 GPUs.